In a crank and slotted lever quick return motion mechanism, the ratio of maximum

velocity at return stroke to the maximum velocity at cutting stroke is 2. If 25cm is the

stroke length, find (a) slotted lever length, (b) the ratio of time of cutting to return

strokes, (c) If the crank rotates at 30rpm, what is the maximum cutting velocity per

sec.


1
Expert's answer
2022-02-23T04:48:01-0500

"(a)\ \dfrac{x}{25} = 2"


"x = 50\ cm"


"(b)\ \sin\dfrac{25}{50}= \sin({90-\dfrac{\theta}2)}"


"30° = 90 -\dfrac{\theta}2"


"\theta= 120°"


ratio of time of cutting to return strokes = "\dfrac{180-\theta}{\theta}" = "0.5"


"(c)\ 30\ rpm × 2 = 60\ rpm"